hiiiihiiiiiiiiii:

Subtract the fractions 5/8-1/8. Remember to reduce. 6/8 1/2 4/0 4/8

justus:

Just subtract the numerators.

justus:

Keep the denominator the same.

hiiiihiiiiiiiiii:

huh

Laylalyssa:

Ok so 5/8-1/8 5+ -1 over 8 =4/8 simplify 4/8 to get ur answer

hiiiihiiiiiiiiii:

so D

justus:

it's not 5 +-1 it's simply 5-1

Laylalyssa:

it said to simplify

justus:

4/8 divide top and bottom by 2

justus:

I mean 4

justus:

4/4= 8/4= ?/? is your simplified answer